Two trains A and B 100m and 60m long are moving in opposite direction on the parallel tracks. The speed of shorter is 3 times the larger one. Of the train take 4 seconds to cross each other then find the velocities of the trains

Grade:11

1 Answers

Khimraj
5 years ago
let the speed of larger train is v
then speed of shorter train is 3v
then relative velocity = 3v – (-v) = 4v
relative displacement = 100 – (-60) = 160
then time taken 
t = 160/4v = 4
40/v = 4
so v = 10 m/s
then
speed of larger train v = 10m/s
speed of shorter train 3v = 30 m/s
Hope it clears. If  you liike answer then please approve it.